# Optimizing Data Transfer Performance and Energy Efficiency with Deep Reinforcement Learning

### v1

@@ Abstract
The rapid growth of data across fields of science and industry has increased the need to improve the performance of end-to-end data transfers while using the resources more efficiently. In this paper, we present a dynamic, multiparameter reinforcement learning (RL) framework that adjusts application-layer transfer settings during data transfers on shared networks. Our method strikes a balance between high throughput and low energy utilization by employing reward signals that focus on both energy efficiency and fairness. The RL agents can pause and resume transfer threads as needed, pausing during heavy network use and resuming when resources are available, to prevent overload and save energy. We evaluate several RL techniques and compare our solution with state-of-the-art methods by measuring computational overhead, adaptability, throughput, and energy consumption. Our experiments show up to 25% increase in throughput and up to 40% reduction in energy usage at the end systems compared to baseline methods, highlighting a fair and energy-efficient way to optimize data transfers in shared network environments.
1 Introduction
The unprecedented increase in data generated by scientific research, industrial applications, e-commerce, social networks, the Internet of Things (IoT), and emerging trends like large-scale Artificial Intelligence (AI) training workloads has created a significant demand for distributed data access, sharing, and dissemination. Consequently, global data traffic has surged exponentially, surpassing an annual rate of 5 zettabytes in 2023 [1]. To put this into perspective, this magnitude of data transfer equates to over 3 billion DVDs being shipped per day throughout the entire year over communication networks.
To meet the requirement for high-speed data movement in distributed and collaborative science projects, high-speed research networks like Internet2 [2] and ESnet [3] have been established. However, despite the availability of these high-speed networks, most data transfers achieve only a fraction of the available bandwidth due to inefficient tuning and underutilization of the available resources. There has been a considerable amount of work on network throughput optimization, proposing heuristic [4, 5, 6, 7, 8], supervised learning [9, 10, 11], and real-time optimization [12, 13, 14, 15] models to address performance limitations. While heuristic models can provide higher transfer throughput than baseline configurations, they often lack robustness in the face of dynamic network conditions. Supervised learning models, when combined with real-time probing, require extensive historical data collection under a wide range of transfer conditions, a process that can span weeks or months. Real-time optimization algorithms, although effective in finding near-optimal settings, typically suffer from long convergence times and may compromise fairness and stability in shared network environments.
Moreover, the exponential increase in data transfers has raised significant concerns over energy consumption, particularly in high-performance computing (HPC) and cloud data centers. It is estimated that information and communication technologies will consume between 8% and 21% of the world’s electricity by 2030 [16], with communication networks accounting for approximately 43% of the total IT power consumption [17]. Data transfers over the Internet now consume more than a hundred terawatt-hours of energy annually, costing around 20 billion US dollars, and recent studies even suggest that, in some scenarios, physically shipping storage media could be less carbon-intensive than network transfers [18]. Despite extensive research in power management techniques for networking infrastructure, little attention has been paid to reducing the energy consumption of the end systems (i.e., sender and receiver nodes) during active data transfers. In fact, one study shows that approximately 25% of the total electricity consumed during end-to-end data transfers occurs at the end systems on a global (intercontinental) network, a figure that increases to 60% on a nationwide network and up to 90% on a local area network [19]. Since this ratio depends on both the number of network devices (e.g., routers, switches) between sender and receiver and the power consumed by each device, decreasing end-system power consumption can lead to significant energy savings. These challenges underline the need for solutions that optimize both throughput and energy efficiency without sacrificing fairness among users in data transfer applications.
To overcome these issues, we introduce a novel deep reinforcement learning (DRL) framework, called SPARTA (Smart Parameter Adaptation via Reinforcement Learning for data Transfer Acceleration), which is built on the following key innovations:
@@ 1 Introduction
  • We train DRL agents using reward signals that are designed to balance energy consumption with fairness. This enables agents to dynamically pause and resume transfer threads based on real-time network conditions, thereby preventing resource oversaturation and reducing energy usage.
@@
  • By intelligently adjusting application-level parameters, the agents can mitigate the risk of overloading network bandwidth and CPU cores during peak times while exploiting available resources during off-peak periods. This dual adaptation not only accelerates data transfers but also maintains fairness among multiple users.
@@
  • To expedite the DRL training process, we introduce an emulation environment that leverages state transition logs captured during initial real-world training episodes. By using transitions from early episodes, our agents can efficiently learn optimal behaviors without the prohibitive time and energy costs associated with extended real data transfers.
@@
By incorporating these innovations, SPARTA is capable of achieving substantial improvements in data transfer performance. Extensive experiments demonstrate up to 25% improvement in throughput and up to 40% reduction in energy consumption compared to baseline methods.
Compared to our previous work [20], which focused solely on tuning TCP stream counts for performance optimization, SPARTA broadens the scope by adopting a multi-parameter DRL framework that simultaneously optimizes concurrency ($cc$) and parallelism ($p$) while incorporating energy efficiency and fairness in the reward function. This comprehensive approach ensures that data transfers are not only faster but also more energy efficient and fair in shared network environments.
@@
The rest of the paper is organized as follows: Section II provides background on the challenges of energy efficiency and fairness in high-performance network data transfers, along with related work. Section III details the design and implementation of our proposed reinforcement learning framework. Section IV presents extensive experimental evaluations of SPARTA, and Section V concludes the paper with insights and future research directions.
2 Background and Related Work
(a) Throughput vs Parameter Settings
(b) Energy vs Parameter Settings
@@ 2 Background and Related Work
Figure 1: The plots show the throughput and energy consumption for different concurrency and parallelism settings under varying background traffic conditions observed at different times of the day in the Chameleon Cloud between TACC and UC sites, connected by a 10 Gbps network link. Subfigure (a) highlights the throughput behavior, while subfigure (b) shows the average energy consumption per measurement interval (MI) of 1 second $(power)$, both influenced by the chosen levels of concurrency and parallelism. These results demonstrate the performance and energy efficiency trade-offs under different background traffic.
There is a considerable amount of work in the literature that tries to improve the performance of data transfers at different layers. Some work focused on developing new transport layer protocols to maximize bandwidth utilization  [21, 22, 23, 24]. However, the wide adoption of a new transport layer protocol requires kernel and system modification, making this type of solution challenging. Additionally, file transfers in high-speed networks frequently encounter I/O performance constraints, enhancing congestion control algorithms alone proves insufficient to address the performance challenges present in contemporary high-performance networks.
An effective strategy to mitigate performance issues in file transfers involves adjusting application-layer transfer parameters, including pipelining ($pp$), parallelism ($p$), concurrency ($cc$), buffer size and striping  [13]. Among these parameters, parallelism and concurrency are identified as the most critical for resolving the majority of underlying performance challenges [4, 13]. Concurrency (cc) refers to task-level parallelism, where each task is the transfer of a file. By running multiple server processes or threads simultaneously, different files can be transferred at the same time, speeding up the overall transfer process. Parallelism (p) denotes the number of parallel streams for each file transfer process or thread. It is beneficial for transferring medium sized and large files. Increasing the total number of data streams (cc $\times$ p) can enhance the transfer throughput. However, excessive use can lead to network congestion and trigger TCP’s congestion avoidance algorithm, reducing the sending rate. When meticulously optimized, these parameters can substantially enhance end-to-end data transfer performance by addressing common bottlenecks such as inadequate I/O parallelism and TCP buffer size limitations  [20].
@@
Previous research has explored various methods for tuning application-layer parameters, focusing on maximizing throughput. Heuristic-based systems [4, 5, 6, 7, 8] often outperform simple single-stream baselines but do not adapt well to changing network conditions (e.g., varying background traffic in different networks). Supervised learning approaches [9, 10, 11] leverage historical data to predict optimal parameter settings. However, gathering large, representative datasets under diverse transfer conditions can be time-intensive and expensive. Online methods [12, 13, 15] continuously adjust parameters but may experience lengthy convergence times and lack fairness safeguards in multi-tenant environments.
@@
A popular data transfer solution, Globus [25], uses static concurrency (2–8) and does not respond to changing network conditions, often underutilizing available bandwidth. ProbData [26] applies stochastic approximation to tune parameters but can require hours to converge, while PCP [13] relies on hill-climbing in a reduced search space, limiting speed and precision. Some earlier work [4, 11] employed heuristic and historical-data models to estimate parameter settings that maximize throughput; however, such solutions depend on extensive historical logs from production networks. In many cases, the cost and complexity of assembling these logs renders them impractical.
There has been also some prior work on energy-efficient data transfer optimization. Most existing work in this area has centered on reducing power usage within the network core (e.g., switches and routers) [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. While these approaches reduce overall network energy consumption, they can also degrade performance (e.g., by putting idle components to sleep) or require costly hardware replacements, limiting their practicality in the short term.
@@
By comparison, fewer studies address energy use in end-to-end data transfers, particularly emissions at the sender and receiver systems during active transfers. Existing methods—ranging from heuristic-based [42, 5, 43] to historical-data-driven [10, 9]—typically focus on adjusting transfer parameters while honoring throughput or energy targets specified by service-level agreements (SLAs). However, building robust historical models demands extensive data covering diverse traffic conditions, which is both time-consuming and costly, and heuristics tend to falter when applied to new or different network settings.
3 Model Overview
3.1 Why Reinforcement Learning?
@@
Existing heuristic solutions cannot easily adapt to the dynamic nature of wide-area networks, where optimal concurrency ($cc$) and parallelism ($p$) depend on both static factors (e.g., hardware limits, available network bandwidth) and highly variable conditions (e.g., background traffic, transient congestion). Figure 1 displays throughput and energy consumption for various concurrency and parallelism settings under different network traffic conditions in the Chameleon Cloud [44] between TACC and UC sites. These sites are connected by a 10 Gbps network, and the experiment involved transferring 50 files of 1 GB each using TCP CUBIC [45]. The plots illustrate how varying the levels of $cc$ and $p$ influences receiver-side throughput and energy usage. The results highlight the trade-offs between performance and energy efficiency under various network conditions. The relationship between the parameters is non-linear, and the optimal settings can improve performance by up to 10 times compared to baseline settings (where concurrency ($cc$) = 1 and parallelism ($p$) = 1). As background traffic changes, the optimal settings for throughput and energy also shift. With limited network signals available from end hosts, an ideal data transfer solution needs to adjust $cc$ and $p$ to maintain optimal performance and energy consumption.
To understand why an adaptable solution is required, we consider the well-known models for single- and multi-stream TCP throughput. In this study, we analyze loss-based TCP variants, such as TCP CUBIC [45], for all our analysis. When the packet loss ratio $L<1\%$, a single TCP flow’s throughput can be approximated as [46]:
$TCP_{thr}\leq\frac{MSS}{RTT}\cdot\frac{C}{\sqrt{L}},$ (1)
where $MSS$ is the maximum segment size, $RTT$ is the round-trip time, $C$ is a constant factor, and $L$ is the packet loss ratio. For $n$ parallel TCP streams, the aggregate throughput can be approximated by summing each flow’s bandwidth [47]:
$TCP_{agg}\leq\frac{C}{RTT}\Bigl{[}\frac{MSS}{\sqrt{L_{1}}}+\dots+\frac{MSS}{\sqrt{L_{n}}}\Bigr{]}.$ (2)
While adding parallel streams can boost throughput by effectively enlarging a “virtual” $MSS$, the benefits vanish if packet loss rises from congestion or if the bottleneck link is fully saturated. Consequently, an optimal stream count exists that maximizes throughput without triggering excess loss. However, due to fluctuating background traffic, this “optimal” stream count varies throughout the transfer session and across different network paths.
Heuristic methods for fixed or slowly adapting concurrency and parallelism struggle in dynamic environments. We need a framework that: (1) requires no prior knowledge of the network or endpoints, (2) derives insights solely from local observations (e.g., throughput, RTT, packet loss) available at end hosts, (3) learns the mapping between these signals and the current network congestion state, and (4) handles partial observability (i.e., incomplete visibility of network internals).
Reinforcement Learning (RL), especially Deep Reinforcement Learning (DRL), fulfills these criteria by treating concurrency ($cc$) and and parallelism ($p$) parameter selection as a sequential decision-making problem. The agent observes performance signals (e.g., RTT and its derivatives, packet loss rate), takes an action (e.g., increasing/decreasing $cc$ or $p$), and receives a reward that encodes its objectives (e.g., higher throughput, lower energy usage, or fairness).
Because network congestion states are not fully observable from the end host, the agent’s policy function approximator infers hidden conditions by continuously monitoring above mentioned feedback signals as state. Over time, the DRL algorithm discovers the “sweet spot” for concurrency and parallelism level at any given instant, dynamically reacting to background traffic changes.
3.2 Optimization Goals and Problem Formulation
@@
Our novel solution, SPARTA, provides a flexible way to optimize different objectives by adjusting the reward function during training. It aims to maximize data throughput while either (1) promoting fairness and indirectly reducing energy via low congestion, or (2) prioritizing throughput per unit energy.
3.2.1 Fairness and Efficiency Objective
@@
A key aim in shared networks is to achieve high throughput without disproportionately harming other transfers sharing the link. We adopt a utility function $U(T,L)$ [15] that rewards higher throughput $T$ and penalizes loss $L$:
$U(T,L)=\frac{T}{K^{(cc\times p)}}-T\cdot L\cdot B,$ (3)
@@
where $K$ and $B$ are constants controlling the relative weight of throughput scaling and the cost of loss. By discouraging congestion (i.e., excessive packet loss), this objective not only ensures fairness but also reduces retransmissions and idle waiting, indirectly improving energy efficiency.
The goal is to choose concurrency ($cc$) and parallelism ($p$) that maximize the cumulative utility between transfer start time $t_{s}$ and finish time $t_{f}$:
$\underset{\{cc,p\}}{\operatorname{argmax}}$ $\quad\sum_{t=t_{s}}^{t_{f}}U(T_{t},L_{t})$ (4)
subject to: $\quad cc\times p\leq\mathbb{N}_{\text{streams}},$
$\quad T_{t}\leq b,$
where $T_{t}$ and $L_{t}$ are the throughput and packet loss rate at interval $t$, and $b$ is the link capacity. By avoiding congestion, the agent prevents wasted energy from retransmissions and queuing delays, thereby promoting both fairness and efficiency.
3.2.2 Throughput-Focused Energy Objective
Some applications prioritize minimizing energy consumption while still achieving acceptable throughput. To address this, we use the throughput-to-energy ratio $\tfrac{T_{t}}{E_{t}}$ at each interval $t$, where $T_{t}$ is the throughput and $E_{t}$ is the energy consumption:
$\underset{\{cc,p\}}{\operatorname{argmax}}$ $\sum_{t=t_{s}}^{t_{f}}\frac{T_{t}}{E_{t}}$ (5)
subject to: $\quad cc\times p\leq\mathbb{N}_{\text{streams}},$
$\quad T_{t}\leq b,$
where $b$ is the available bandwidth. This formulation explicitly rewards higher throughput per unit of energy, making it especially relevant in energy-constrained environments.
3.2.3 Adaptive DRL Framework and Interrelation of Objectives
Both the fairness-and-efficiency (Eq. 4) and throughput-focused energy (Eq. 5) objectives can be integrated into the same DRL framework via a customizable reward:

### v2

# Energy-Efficient and High-Performance Data Transfers with DRL Agents

@@ Abstract
The rapid growth of data across fields of science and industry has increased the need to improve the performance of end-to-end data transfers while using the resources more efficiently. In this paper, we present a dynamic, multiparameter deep reinforcement learning (DRL) framework that adjusts application-layer transfer settings during data transfers on shared networks. Our method strikes a balance between high throughput and low energy utilization by employing reward signals that focus on both energy efficiency and fairness. The DRL agents can pause and resume transfer threads as needed, pausing during heavy network use and resuming when resources are available, to prevent overload and save energy. We evaluate several DRL techniques and compare our solution with state-of-the-art methods by measuring computational overhead, adaptability, throughput, and energy consumption. Our experiments show up to 25% increase in throughput and up to 40% reduction in energy usage at the end systems compared to baseline methods, highlighting a fair and energy-efficient way to optimize data transfers in shared network environments.
Index Terms:
Green computing; sustainability; energy efficient data transfers; application-layer optimization; deep reinforcement learning.
I Introduction
The explosive growth of data generated by scientific research, industrial applications, e-commerce, social networks, the Internet of Things (IoT), and large-scale AI training workloads has driven global data traffic past 5 zettabytes per year as of 2023 [1]—equivalent to shipping over 3 billion DVDs daily. Accessing, sharing, and disseminating data at this scale in an efficient and sustainable manner remains an open challenge.
A critical dimension of this challenge is energy consumption. Information and communication technologies are projected to consume between 8% and 21% of global electricity by 2030 [12], with communication networks alone accounting for roughly 43% of total IT power draw [6]. Internet data transfers already consume over a hundred terawatt-hours annually, costing approximately 20 billion US dollars [47], and network energy costs already represent 40–50% of operational expenditures for telecommunications operators in developing countries [40]. Recent analyses warn that the cost of data movement already dominates electrical losses and may undermine future gains in compute energy efficiency if left unaddressed [54]. In some scenarios, physically shipping storage media is less carbon-intensive than network transfer [10].
Despite extensive work on power management for networking infrastructure, energy consumption at the end systems (sender and receiver nodes) during active transfers has received little attention. Yet end-system power can account for approximately 25% of total transfer energy on intercontinental paths, rising to 60% on nationwide networks and up to 90% on local area networks [3]. Because this ratio depends on the number of intermediate network devices and their per-device power draw, reducing end-system consumption can yield substantial savings.
Achieving these savings requires parameter adaptation strategies that cope with highly dynamic conditions—fluctuating bandwidth, varying round-trip times, and competing cross-traffic. Traditional heuristic approaches lack the robustness to generalize across such diverse scenarios. To address this, we propose SPARTA (Smart Parameter Adaptation via Reinforcement Learning for data Transfer Acceleration), a deep reinforcement learning (DRL) framework with the following contributions:
@@ I Introduction
  • We train DRL agents with reward signals that balance energy consumption against fairness, enabling agents to dynamically pause and resume transfer threads based on real-time conditions to prevent resource oversaturation and reduce energy usage.
@@
  • By jointly adjusting transfer parameters, agents avoid overloading bandwidth and CPU cores during peak periods while exploiting available resources during off-peak periods, accelerating transfers while preserving fairness among concurrent users.
@@
  • We introduce an emulation environment that replays state-transition logs from early real-world episodes, enabling efficient policy learning without the prohibitive time and energy costs of prolonged live transfers.
@@
  • Experiments demonstrate up to 25% throughput improvement and up to 40% reduction in energy consumption compared to baseline methods.
@@
The remainder of the paper is organized as follows: Section II provides background and related work on energy-efficient high-performance data transfers. Section III details the design and implementation of the proposed framework. Section IV presents experimental evaluations, and Section V concludes with insights and future directions.
II Background and Related Work
@@
Figure 1: The plots show the throughput and energy consumption for different concurrency and parallelism settings under varying background traffic conditions observed at different times of the day in the Chameleon Cloud between TACC and UC sites, connected by a 10 Gbps network link. Subfigure (a) highlights the throughput behavior, while subfigure (b) shows the energy consumption both influenced by the chosen levels of concurrency and parallelism. These results demonstrate the performance and energy efficiency trade-offs under different background traffic
II-A Application-Layer Parameter Tuning
Improving data transfer performance at the transport layer—e.g., via new congestion control protocols [56, 37, 14, 17]—requires kernel and system modifications that hinder wide adoption. A more practical strategy is tuning application-layer parameters such as pipelining ($pp$), parallelism ($p$), concurrency ($cc$), buffer size, and striping [58], of which parallelism and concurrency are the most impactful [11, 58, 34].
Concurrency ($cc$) is task-level parallelism: multiple server processes or threads transfer different files simultaneously. Parallelism ($p$) is the number of TCP streams per file transfer, primarily beneficial for medium and large files. Increasing the total stream count ($cc\times p$) improves throughput, but excessive streams provoke TCP congestion avoidance and reduce the sending rate. When properly tuned, these parameters address common bottlenecks such as inadequate I/O parallelism and TCP buffer limitations [30].
II-B Throughput Optimization
@@
Prior approaches to application-layer tuning fall into three categories. Heuristic-based systems [9, 11, 16, 43, 27, 49, 36, 8] outperform single-stream baselines but generalize poorly to dynamic network conditions. Supervised learning methods [29, 52, 46] predict optimal settings from historical data, yet assembling large, representative datasets under diverse conditions is expensive. Online methods [41, 58, 7] adapt continuously but suffer from slow convergence and lack fairness safeguards in multi-tenant environments.
@@
Widely used tools reflect these limitations. Globus [4] applies static concurrency (2–8) and cannot react to changing conditions, often underutilizing available bandwidth. ProbData [59] uses stochastic approximation but can require hours to converge, while PCP [58] employs hill-climbing over a reduced search space, limiting both speed and precision. Other work [11, 46] relies on heuristic or historical-data models whose dependence on extensive production-network logs makes them impractical in many settings.
II-C Energy Optimization
Most energy-efficiency research targets the network core—switches and routers [13, 50, 60, 22, 15, 18, 19, 51, 35, 25, 57, 5, 42, 20, 48]. These techniques reduce network-level consumption but can degrade performance (e.g., by sleeping idle components) or require costly hardware upgrades, limiting near-term practicality.
@@ II-C Energy Optimization
Far fewer studies address energy consumption at the end systems during active transfers. Existing approaches—heuristic-based [2, 16, 21] and historical-data-driven [52, 29]—adjust transfer parameters to meet throughput or energy targets specified by SLAs. However, building robust historical models demands extensive data across diverse traffic conditions, and heuristics generalize poorly to unseen network settings.
In our prior work [30], we tune TCP stream counts for throughput optimization. SPARTA extends this by adopting a multi-parameter DRL framework that jointly optimizes concurrency ($cc$) and parallelism ($p$) while incorporating energy efficiency and fairness directly into the reward function, ensuring transfers that are faster, more energy-efficient, and fair in shared network environments.
III SPARTA Model Overview
In SPARTA, we use Reinforcement Learning (RL) effectively for balancing the energy consumption during data transfers with the achieved throughput and fairness. In the following subsections, we discuss why we choose RL and the details of our RL solution.
III-A Why Reinforcement Learning?
@@
Existing heuristic solutions cannot easily adapt to the dynamic nature of wide-area networks, where optimal concurrency ($cc$) and parallelism ($p$) depend on both static factors (e.g., hardware limits, available network bandwidth) and highly variable conditions (e.g., background traffic, transient congestion). Figure 1 displays throughput and energy consumption for various concurrency and parallelism settings under different network traffic conditions in the Chameleon Cloud [33] between TACC and UC sites. These sites are connected by a 10 Gbps network, and the experiment involved transferring 100 files of 1 GB each using TCP CUBIC [23]. The plots illustrate how varying the levels of $cc$ and $p$ influences receiver-side throughput and energy usage. The results highlight the trade-offs between performance and energy efficiency under various network conditions. The relationship between the parameters is non-linear, and the optimal settings can improve performance by up to 10 times compared to baseline settings (where concurrency ($cc$) = 1 and parallelism ($p$) = 1). As background traffic changes, the optimal settings for throughput and energy also shift. With limited network signals available from end hosts, an ideal data transfer solution needs to adjust $cc$ and $p$ to maintain optimal performance and energy consumption.
To motivate the need for adaptive control, we recall classic loss-based TCP throughput behavior (TCP CUBIC in our experiments). When packet loss ratio $L<1\%$, a single TCP flow throughput can be approximated [44] as:
$TCP_{\mathrm{thr}}\lesssim\frac{MSS}{RTT}\cdot\frac{C}{\sqrt{L}},$ (1)
and for $n$ parallel TCP streams, the aggregate throughput can be approximated by summing per-flow rates [24] as:
$TCP_{\mathrm{agg}}\lesssim\frac{C}{RTT}\sum_{i=1}^{n}\frac{MSS}{\sqrt{L_{i}}}.$ (2)
Increasing streams can increase aggregate throughput until the bottleneck saturates. Beyond that point, congestion increases loss and round trip time (RTT), which reduces goodput and increases energy waste (retransmissions and waiting). Because the bottleneck state changes over time, heuristics with fixed or slowly adapting parameters are not ideal.
We therefore model this adaptive control as a sequential decision problem and use Deep Reinforcement Learning (DRL) to learn a policy that maps end-host signals to parameter updates, without requiring prior knowledge of network internals.
III-B Optimization Goals and Formulation
At each monitoring interval (MI) $t$, the controller chooses concurrency and parallelism $(cc_{t},p_{t})$ and creates a TCP stream budget:
$n_{t}\triangleq cc_{t}\cdot p_{t},\qquad n_{t}\in\{1,2,\dots,N\},$ (3)
where $N$ is a safe number of TCP streams denoting end-host and network stability constraints.
@@ III-B Optimization Goals and Formulation
Let $T_{t}$, $L_{t}$, and $E_{t}$ be the measured throughput, packet loss rate, and end-system energy during MI $t$. SPARTA aims to maximize data throughput while (1) promoting fairness and indirectly reducing energy via low congestion, or (2) prioritizing throughput per unit energy.
(a) Utility (Fairness & Efficiency) Surface over $(cc,p)$
(b) Utility (Throughput/Energy) Surface over $(cc,p)$
Figure 2: Utility geometry and trackability. (a) and (b) Utility induced by Eq. (4) and Eq. (6) over the feasible $(cc,p)$ grid exhibits a single dominant peak near the knee region, consistent with a unimodal/quasi-concave profile in $n=cc\cdot p$ for a fixed network condition.
III-B1 Fairness and Efficiency Objective (SPARTA-FE)
@@
In shared networks, high throughput should not be achieved by pushing the bottleneck into persistent congestion that harms other transfers. We adopt a utility function  [7] that rewards throughput while penalizing congestion (loss):
$U_{\mathrm{FE}}(T_{t},L_{t},n_{t})\triangleq\frac{T_{t}}{K^{n_{t}}}-B\,T_{t}\,L_{t},$ (4)
@@
where $K>1$ and $B>0$ are constants controlling the relative weight of throughput scaling and the cost of loss. By discouraging congestion (i.e., excessive packet loss), this objective not only ensures fairness but also reduces retransmissions and idle waiting, indirectly improving energy efficiency.
Over a transfer session $t_{s},t_{f}]$, the objective is to maximize cumulative utility:
$\max_{\{cc_{t},p_{t}\}}\sum_{t=t_{s}}^{t_{f}}U_{\mathrm{FE}}(T_{t},L_{t},n_{t})\quad\text{s.t.}\quad 1\leq n_{t}\leq N,\;T_{t}\leq b,$ (5)
where $b$ is the upper bound on achievable throughput.
III-B2 Throughput-Focused Energy Objective (SPARTA-T)
For energy-sensitive deployments, we optimize throughput per unit energy:
$U_{\mathrm{TE}}(T_{t},E_{t})\triangleq\frac{T_{t}}{E_{t}},$ (6)
and maximize
$\max_{\{cc_{t},p_{t}\}}\sum_{t=t_{s}}^{t_{f}}U_{\mathrm{TE}}(T_{t},E_{t})\quad\text{s.t.}\quad 1\leq n_{t}\leq N,\;T_{t}\leq b.$ (7)
III-B3 Unifying View
Both objectives are implemented through a selectable reward: